Giving the following data for XY Company

year

2008

2009

Total Sales

$145,0000

$121,4000

Cost of goods sold

$880,000

$738,400

Selling & Administrative Expenses

$92,500

$80,700

Assume high-low method is adequate in this case, the variable costs of cost of goods sold per unit is

1.

40%

2.

50%

3.

60%

4.

65%
Answer #1

Answer-----------(3) 60%

Working

Sales $    1,214,000.00
Variable cost (1214000 x 0.60) $       728,400.00
Fixed cost $          10,000.00
Total cost $       738,400.00
Variable cost ratio (728400/1214000) 60%

Working note 2

For this question i have assumed that sales price is $1 per unit so it gives sales units as equal to sales revenue.

Cost No. of activities
A High Level $        880,000.00 1450000
B Low Level $        738,400.00 1214000
C=A-B Difference $        141,600.00                    236,000
A Cost difference $        141,600.00
B No. of activities difference 236000
C=A/B Variable cost per unit $                     0.60
High Level Low Level
A No. of activities 1450000 1214000
B Variable cost per unit $                0.6000 $                  0.6000
C=AxB Total Variable cost $        870,000.00 $         728,400.00
D Total cost $        880,000.00 $         738,400.00
E=D-C Total fixed cost $          10,000.00 $            10,000.00